OpenStudy (anonymous):

Please Help!! Find the surface area of the curved portion of the beer keg obtained be revolving the curve x=sqr(16-y^2) for -3≤y≤3 about the Y axis. Omit the area of the flat top and bottom portions.

OpenStudy (anonymous):

The formula for the area of the surface obtained by revolving x=f(y) from y=a to y=b about the y-axis is \[2\pi \int\limits_{a}^{b}f(y)\sqrt{1+(f'(y))^2}dy\]